# Evaluation of Language Models in the Medical Context Under Resource-Constrained Settings

### v1

@@ 5.5 The groups of generative models: LLaMA and GPT
Two procedures are carried out to gain insight into whether the models exhibit similar perplexity behavior and to identify potential clusters among them. The first procedure involves calculating the correlations between the models. Spearman’s and Pearson’s correlations are considered, assessing monotonic and linear relationships. The second procedure involves dimensionality reduction via UMAP, followed by hierarchical clustering, represented in Fig. 6 using dendrograms.
The correlation analysis reveals that, in general, all models are positively related. This positive relation indicates that most samples have a similar relative difficulty level for the models. However, the BioGPT models are the only exceptions, showing low Spearman correlations compared to the other models and positioning them considerably far from the rest of the evaluated models. The observed behavior in terms of monotonic relationships remains consistent regardless of the use of the BOS token. Regarding linear relationships, groups of models are clearly observed, including LLaMA models, OpenLLaMA models, and GPT-2 models. Although the previous groups are somewhat expected, some unexpected associations are also evident, such as between Falcon-7B and MPT-7B and between Palmyra Base 5B and GPT-J 6B. These two groups have notably weak intergroup correlations.
With respect to the LLaMA and OpenLLaMA models, their linear relationship weakens when the BOS token is used, indicating more pronounced performance disparities. Possibly, training data plays an important role in this context, as it is essentially their main difference [92]. Further insights, provided by the correlations and supported by the dendrograms in Fig. 6, reveal the existence of two main groups of models: the GPT-2 models and the LLaMA models.
The analysis of the standard deviations reveals the presence of exceptionally challenging samples for the models, that is, outliers. This phenomenon is visually depicted in Fig. 4, which illustrates the distribution of perplexity between different models. Overall, the perplexity distributions have considerable variance, with the LLaMA models demonstrating the slightest variation. More detailed analyses are required to estimate the importance of these outliers and possible patterns among them.
@@ 6 Conclusion
This study comprehensively explored a range of language models with varying sizes and domain fields. These models, encompassing 52 in total, were tested in two fundamental natural language processing tasks: text classification and text generation. The size range of these models goes from 110 million to 13 billion parameters, which is relatively small compared to recent language models. Our findings, based on these specific models, have significant implications, particularly for researchers and organizations operating under resource-constrained contexts.
@@
In text classification, our research has explored the effectiveness of three distinct approaches: context embedding similarity, natural language inference, and multiple-choice question-answering. BioLORD and SapBERT models have demonstrated remarkable performance in text classification via contextual embedding similarity. Additionally, the instruction-tuned versions of T5, Flan-T5 and T0, followed by the instruction-tuned versions of LLaMA, have exhibited good results in the multiple-choice question-answering approach. However, for the NLI approach, we have identified the need for extended exploration, particularly in the context of specialized domain models. This implies that further research is needed to fully understand the potential of NLI in text classification, especially in specialized domains.
@@
A common thread running through our findings is the significance of the prompt in improving text classification performance across different datasets and approaches. The significance of prompts extends beyond mere performance gains; they present a viable alternative to the resource-intensive processes of training and fine-tuning language models, which are often associated with substantial financial and environmental costs, especially for large-scale models. Moreover, effective prompt engineering is essential to mitigate prompt brittleness, ensuring more robust and reliable outcomes.
@@
Since, in general, medical datasets remain relatively small and cover only a small region of the medical knowledge space [79], domain-specific models, specialized by continuous pre-training on only one of these datasets, might see their generalization ability hindered. This practice could be one of the explanations for the results obtained. The results also suggest that the architecture, training data, and training objectives are crucial in determining the model’s generalization abilities, possibly outweighing model size.
@@
Regarding the text generation task, two groups of models are identified based on the perplexities obtained on MIMIC-CXR. These groups consist of the GPT-2 models and the LLaMA models. Notably, the LLaMA models stand out due to their low perplexities with minimal variation. Further research is needed to identify and understand the outliers within these results, as they could hold important insights.
@@
In conclusion, this research highlights the critical role of prompts in language model inference and reaffirms the effectiveness of instruction-tuned generative models in addressing downstream tasks. It also highlights the significance of model architecture, training data, and training objectives, potentially even more so than model size, in its generalization capacity. We advocate for further investigations into topics such as model calibration, i.e., how certain the model is about output, prompt tuning, and performance concerning issues like hallucinations and biases, among others. Such studies can lead to more effective and ethical applications of language models, particularly in sensitive fields like healthcare.
@@ A.3.2 Description
Appendix B Prompts
@@
B.1 Prompts for text classification via contextual embedding similarity and NLI
The prompts proposed for text classification using contextual embedding similarity and Natural Language Inference (NLI) are exclusively applied to the label (in the case of NLI, to the hypothesis). Table 3 lists the prompts used. Template ID 0 is the default to generate the hypothesis in the zero-shot text classification using the NLI setting, as documented in HuggingFace.
| ID | Label prompt | Dataset |
@@ B.1 Prompts for text classification via contextual embedding similarity and NLI
B.2 Prompts for text classification via multiple-choice QA
The proposed prompts for text classification via multiple-choice question answering are based on the default templates specific to various of the considered instruction-tuned models. These templates are systematically assessed using a set of questions, enabling us to quantify the influence of the question wording. For the MS-CXR dataset, we also incorporate role-based questions. The prompts, their corresponding datasets, and specific requirements are summarized in Table 5 and Table 4.
@@ B.2 Prompts for text classification via multiple-choice QA
| ID | Prompt | Requirements | Dataset |
@@
Appendix C Supplementary results
This appendix presents supplementary figures and tables to support the results presented. They are displayed first by dataset and then by task or approach.
C.1 Classification task
C.1.1 Contextual embedding similarity
@@
C.1.2 Natural language inference
@@
C.1.3 Multiple choice question answering

### v2

@@ 5.2 Conditional text generation analysis
Two procedures are carried out to determine whether the models exhibit similar perplexity behavior and identify potential clusters among them. The first procedure involves calculating the correlations between the models. Spearman’s and Pearson correlations are considered, assessing monotonic and linear relationships, respectively. The second procedure consists of dimensionality reduction via UMAP, followed by hierarchical clustering, represented by dendrograms in Fig. 6. Both procedures reveal the existence of two main groups of models: the GPT-2 and the LLaMA models.
In general, all models are positively correlated, indicating that most samples have a similar relative difficulty for these models. BioGPT models (m48-m49) are the only exception to this. Further looking at the Pearson correlations, clustering patterns are present, where groups such as the LLaMA, the OpenLLaMA, and the GPT-2 models are identified. Although these previous clusters are somewhat expected, some unexpected associations are also evident, such as between Falcon-7B and MPT-7B and between Palmyra Base 5B and GPT-J 6B. Moreover, linear relationships between the LLaMA and OpenLLaMA models weaken, interestingly, when the BOS token is used, indicating more pronounced performance disparities.Possibly, training data plays a role, as it is essentially their main difference [98].
@@ 6 Conclusion
This study comprehensively explores small pre-trained language models with varying sizes, architectural families, and domains. These models, being considered, are tested for two fundamental medical natural language processing tasks: text classification and conditional text generation. The size of the models ranges from million to billion parameters, which is relatively small compared to recent language models but suitable for consumer-grade computing resources. Our findings have significant implications, particularly for researchers and organizations operating under computational resource-constrained settings.
@@
For the text classification task, three distinct approaches are explored: context embedding similarity, natural language inference, and multiple-choice question answering. BioLORD and SapBERT models have demonstrated remarkable performance in text classification via contextual embedding similarity. Similarly, the instruction-tuned versions of T5, Flan-T5 and T0, followed by the instruction-tuned versions of LLaMA, have exhibited outstanding results in the multiple-choice question answering approach. Flan-T5 and T0 are remarkably good in both general medical and radiology-specific knowledge assessments. To fully understand NLI models’ potential, further exploration of this approach is needed, particularly in specialized domains.
@@
A common thread running through our findings is the significance of the prompt in improving text classification performance across different datasets and approaches. This significance extends beyond performance gains; they present a viable alternative to the resource-intensive processes of training and fine-tuning language models, which are often associated with substantial financial and environmental costs. Effective prompt engineering is also essential to mitigate prompt brittleness, ensuring more robust and reliable outcomes. As prompt brittleness is evidenced during the study, and given its importance, further exploration in this line of research is recommended.
@@
Medical datasets often remain relatively small and cover only a small region of the medical knowledge space [84], so domain-specific models specialized using these datasets might see their generalization ability hindered. This practice could explain, to some extent, the results obtained. The results also suggest that the architecture, training data, and training objectives are crucial in determining the model’s generalization abilities, possibly outweighing the relevance of model size as a single variable.
@@
For the conditional text generation task, LLaMA models stand out due to their low perplexities with minimal variation. Two groups of models are also identified based on the perplexities obtained in MIMIC-CXR: a group consisting of GPT-2 models and another of LLaMA models. Further research is needed to identify and understand the outliers within these results, as they could hold important insights.
@@
In conclusion, this research highlights the critical role of prompts in language model inference and reaffirms the effectiveness of instruction-tuned generative models in addressing downstream tasks. It also underscores the relevance of model architecture, training data, and training objectives, potentially even more so than model size alone, in its generalization capacity. We advocate for further investigations into topics such as model calibration, i.e., how certain the model is about output, prompt engineering and tuning, and performance concerning issues like hallucinations and biases, among others. Such studies can lead to more effective and ethical applications of language models in healthcare. Extensions to include quantized models and more medical NLP tasks will be considered in further research. Quantification is an interesting and promising approach to making LLMs viable in consumer-grade computing resources.
@@ A.3.2 Description
Appendix B Models
| ID | Model | Type | Domain | Model size (no. parameters) | Input max. size (no. tokens) | Total inference time (seconds) | |||
| m00 | BERT [16] | Encode-only | General | M | – | ||||
| m01 | BERT [16] | Encode-only | General | M | – | ||||
| m02 |
Encode-only | Biomedical | M | – | ||||
| m03 |
Encode-only | Biomedical | M | – | ||||
| m04 |
Encode-only | Biomedical | M | – | ||||
| m05 | SciBERT [65] | Encode-only | Biomedical | M | |||||
| m06 | SapBERT [90] | Encode-only | Biomedical | M | – | ||||
| m07 | BioLORD-STAMB2-v1 [91] | Encode-only | Biomedical | M | – | ||||
| m08 | BioLORD-STAMB2-v1-STS2 [91] | Encode-only | Biomedical | M | – | ||||
| m09 | BioLORD-PMB [91] | Encode-only | Biomedical | M | – | ||||
| m10 | Bio+Clinical BERT [70] | Encode-only | Clinical | M | – | ||||
| m11 | NLI-DeBERTa [88] |
General | M | – | ||||
| m12 | RoBERTa-MNLI [89] |
General | M | – | ||||
| m13 | BART Large-MNLI [94] | Encoder-decoder | General | M | – | ||||
| m14 | T5-V1.1-Base [92, 93] | Encoder-decoder | General | M | – | ||||
| m15 | T5-V1.1-Large [92, 93] | Encoder-decoder | General | M | – | ||||
| m16 | T5-V1.1-3B [92, 93] | Encoder-decoder | General | B | – | ||||
| m17 | T5-V1.1-11B [92, 93] | Encoder-decoder | General | B | – | ||||
| m18 | Flan-T5-Base [17] |
General | M | – | ||||
| m19 | Flan-T5-Large [17] |
General | M | – | ||||
| m20 | Flan-T5-XL [17] |
General | B | – | ||||
| m21 | Flan-T5-XLL [17] |
General | B | – | ||||
| m22 | T0 3B [95] |
General | B | – | ||||
| m23 | T0++ [95] |
General | B | – | ||||
| m24 | ClinicalT5-base [96] | Encoder-decoder | Clinical | M | – | ||||
| m25 | ClinicalT5-large [96] | Encoder-decoder | Clinical | M | – | ||||
| m26 | GPT-2 Medium [19] | Decoder-only | General | M | – | ||||
| m27 | GPT-2 Large [19] | Decoder-only | General | M | – | ||||
| m28 | GPT-2 XL [19] | Decoder-only | General | B | – | ||||
| m29 | Palmyra Base 5B [97] | Decoder-only | General | B | |||||
| m30 | Camel 5B [99] |
General | B | – | ||||
| m31 | GPT-J 6B [100] | Decoder-only | General | 6.0 B | 2 048 | ||||
| m32 | Instruct GPT-J [101] |
General | B | – | ||||
| m33 | Falcon-7B [102] | Decoder-only | General | B | |||||
| m34 | Falcon-7B-Instruct [102] |
General | B | – | ||||
| m35 | MPT-7B [103] | Decoder-only | General | B | |||||
| m36 | MPT-7B-Instruct [103] |
General | B | – | ||||
| m37 | LLaMA-7B [27] | Decoder-only | General | B | |||||
| m38 | LLaMA 2-7B [28] | Decoder-only | General | B | |||||
| m39 | Alpaca 7B [104] |
General | B | – | ||||
| m40 | LLaMA 2-CHAT-7B [28] |
General | B | – | ||||
| m41 | OpenLLaMA 3B [98] | Decoder-only | General | 3.0 B | 2 048 | – | |||
| m42 | OpenLLaMA 3Bv2 [98] | Decoder-only | General | 3.0 B | 2 048 | – | |||
| m43 | OpenLLaMA 7B [98] | Decoder-only | General | 7.0 B | 2 048 | – | |||
| m44 | OpenLLaMA 7Bv2 [98] | Decoder-only | General | 7.0 B | 2 048 | – | |||
| m45 | OpenLLaMA 13B [98] | Decoder-only | General | 13.0 B | 2 048 | – | |||
| m46 | GPT-2-PubMed Medium [105] | Decoder-only | Biomedical | 355 M | 1 024 | – | |||
| m47 | GPT-2-PubMed Large [105] | Decoder-only | Biomedical | 774 M | 1 024 | – | |||
| m48 | BioGPT [69] | Decoder-only | Biomedical | 347 M | 1 024 | – | |||
| m49 | BioGPT-Large [69] | Decoder-only | Biomedical | 1.5 B | 1 024 | – | |||
| m50 | Galactica 1.3B [26] | Decoder-only | Biomedical | 1.3 B | 2 048 | – | |||
| m51 | Galactica 6.7B [26] | Decoder-only | Biomedical | 6.7 B | 2 048 | – | |||
| m52 | MedAlpaca 7b [71] | Decoder-only | Clinical | B | – | ||||
Appendix C Prompts
@@
C.1 Prompts for text classification via contextual embedding similarity and NLI
The prompts proposed for text classification using contextual embedding similarity and Natural Language Inference (NLI) are exclusively applied to the label (in the case of NLI, to the hypothesis). Table 4 lists the prompts used. Prompt template ID 0 is the default to generate the hypothesis in the zero-shot text classification using the NLI setting, as documented in HuggingFace.
| ID | Prompt template | Dataset |
@@ C.1 Prompts for text classification via contextual embedding similarity and NLI
C.2 Prompts for text classification via multiple-choice QA
The proposed prompts for text classification via multiple-choice question answering are based on the default prompt templates specific to various of the considered instruction-tuned models. These templates are systematically assessed using a set of questions, enabling us to quantify the influence of the question wording. For the MS-CXR dataset, we also incorporate role-based questions. The prompts, their corresponding datasets, and specific requirements are summarized in Table 6 and Table 5.
@@ C.2 Prompts for text classification via multiple-choice QA
| ID | Prompt structure | Requirements | Dataset |
@@
Appendix D Supplementary results
This appendix presents supplementary figures and tables to support the results presented. They are displayed first by dataset and then by task or approach. These results do not are not obtained by bootstrapping, but singular values for the complete inference dataset.
D.1 Text classification task
D.1.1 Contextual embedding similarity
@@
D.1.2 Natural language inference
@@
D.1.3 Multiple choice question answering
@@
D.2 Conditional text generation task
Results are depicted in Fig. 21. The mapping between the models and their ID is
| m26: GPT-2 Medium | m35: MPT-7B | m44: OpenLLaMA 7Bv2 | m50: Galactica 1.3B |
| m27: GPT-2 Large | m37: LLaMA-7B | m45: OpenLLaMA 13B | m51: Galactica 6.7B |
| m28: GPT-2 XL | m38: LLaMA 2-7B | m46: GPT-2-PubMed Medium | |
| m29: Palmyra Base 5B | m42: OpenLLaMA 3B | m47: GPT-2-PubMed Large | |
| m31: GPT-J 6B | m42: OpenLLaMA 3Bv2 | m48: BioGPT-Large | |
| m33: Falcon-7B | m43: OpenLLaMA 7B | m49: BioGPT-Large |